OpenStudy (precal):

Let f(x)=square root (x). If the rate of change of f at x=c is twice its rate of change at x=1, then c=

OpenStudy (precal):

Aren't I looking for the second derivative and just looking for f " (1)?

OpenStudy (anonymous):

no. I understand that as \[f(x)=\sqrt{x}\implies f'(x)={1\over 2\sqrt{x}}\\ f'(x=c)=2\times f'(x=1)\\ {1\over2\sqrt{c}}=2\times{1\over 2\sqrt{1}} \]then solve for 'c'

OpenStudy (anonymous):

rate of change = first derivative

OpenStudy (precal):

no wonder I kept getting -1/4 when the key stated it is +1/4 thanks, I will redo the problem

OpenStudy (precal):

thank you

OpenStudy (anonymous):

you are welcome.